how can I reslove this equation:
60-89(60-89x)/2

Answer by jayanthihemkos@yahoo.com(87)

60-89(60-89x)/2
The sum has to be
60 = 89(60-89x)/2
60*2 = 89(60-89x)/2*2
120 = 89*60 - 89*89x
120 = 5340 - 7921x
120+7921x = 5340-7921x+7921x
7921x+120-120 = 5340 -120
7921x = 5220
7921x/7921 = 5220/7921
x = 5220/7921
x = 0.659